OpenStudy (anonymous):

Write the equation of the line which passes through (–4, 2) and is parallel to y = –x + 6 in slope-intercept form. If you can help then, Thank youuu :)

OpenStudy (anonymous):

y-2=-1(x +4) y-2=-4-x y+x+2=0

OpenStudy (nubeer):

parallel lines have same slope so for your answer slope will be same as the given equation m = -1 points (-4,2) now use this formula y-y1=m(x-x1) just plug in the values.

OpenStudy (anonymous):

m=-1 x=-4 y=2 y=mx+c 2=-1*(-4)+c c=-2 so y=-x-2

OpenStudy (anonymous):

I am sure my answer is right

OpenStudy (anonymous):

Your answer is right! :) Thank you!

OpenStudy (anonymous):

welcome, give me best response

OpenStudy (nurali):

please click best response.